Refer to the following circuit, assuming IS = 1.0 pA; D1 is a generic rectifier diode, and it is at temperature of 500 C. Calculate the voltage drop VR across R1 for


(a) R1 = 1.0 Ω,

(b) R1 = 10 Ω,
(c) R1 = 100 Ω,

(d) R1 = 1.0 kΩ.

R1 1.0Ω D1 V1 -1.ον
Answer #1

Given circuit can be drawn as Below Pg Ri=12 -v-lv VD Current through diode. can be as ID = Is Le hive J-o Where It = 11.600Vox 103 1X2 7.844 1-VD 10 -12 = I x 10 Solving above equation using calculator we we get = Now voltage doop VRI- VVD -0.674 =Note:please don't give dislike .......please comment if you have any problem i will try to solve your problem...... please give thumbs up i am in need....